OpenStudy (anonymous____):

find the missing factor in exponential form 48^6=12^6*__?

jimthompson5910 (jim_thompson5910):

fill in the blank 12 times __________ = 48 tell me what should go in the blank

OpenStudy (anonymous____):

4?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so, \[\Large 48^6 = (12*4)^6\] \[\Large 48^6 = 12^6*4^6\] The missing term is \(\Large 4^6\)

jimthompson5910 (jim_thompson5910):

I'm using the rule that \[\Large (a*b)^c = a^c*b^c\]

OpenStudy (anonymous____):

thank you!

jimthompson5910 (jim_thompson5910):

no problem